# Linear-time classical approximate optimization of cubic-lattice classical spin glasses

### v3

@@ I Introduction
It would seem, then, that one simply needs to find instances of Eq. (1) whose approximate optimization with the best Monte-Carlo algorithms is computationally hard (i.e., exhibits or at least hints at superpolynomial time complexity) in order to identify good candidates for demonstrating time-complexity quantum advantage in heuristic approximate optimization. However, Ref. [17] provides evidence that short-range rugged-energy-landscape spin glasses may be heuristically approximately optimized classically in quadratic time with approximate contractions of tensor networks, though the wall time for cubic lattices with that heuristic proves to be too large to be practical for large systems. Here we present a major advance over that result: We demonstrate a very simple deterministic tensor-network heuristic that is able to compute in linear time (and space) at least some (but not necessarily all) spin configurations that have energies within a few percent of the ground-state energy for the following three versions of Eq. (1): (1) the cubic-lattice model that is used in the D-Wave approximate optimization experiments in Refs. [5] and [6], (2) a planted-solution, cubic-lattice model whose hardness (i.e., time to exact optimization) for state-of-the-art Monte Carlo heuristics is orders of magnitude more than the model, and (3) a cubic-lattice model that encodes (unweighted) Max-Cut on random 3-regular graphs. Unlike the heuristic in Ref. [17], the present one is able to address very large cubic-lattice systems in a practical amount of time. We also provide an explanation for why this heuristic implies that a quantum algorithm must have sublinear time complexity in order to have quantum advantage in heuristic approximate optimization of classical spin glasses that have correlations that decay with euclidean distance, such as models (1) and (2) just mentioned.
@@ V Summary and Discussion
We are hopeful that such a hybridization may provide a new avenue for gaining insights into the properties of classical spin-glass models studied in condensed matter. For example, questions remain open regarding the nature of the low-temperature phase of the cubic-lattice classical model [44]. Monte Carlo methods have been the state-of-the-art approach for this problem, but the hybrid method that we propose would operate according to very different principles and may thereby yield new insights. As another example, Ref. [33] points out that the tunability of the tile-planted-solution model can allow for a systematic study of the interplay between disorder and frustration; the hybrid algorithm that we propose may provide a complementary route to Monte Carlo methods in such a study as well.
We explained that the algorithm may be parallelized through simultaneous use of multiple fragments. This would very substantially reduce the simulation time of our heuristic from what is reported here, which would allow the approximate optimization of much larger systems in a practical amount of time. We also noted that one of the sources of error in our implementation of the algorithm, finite temperature, could be eliminated without changing the time complexity: the ordinary tensors could be replaced by tropical tensors [37]. This would likely reduce the energy errors reported above at the same fragment sizes, but would not be compatible with adapting the algorithm to specialized photonic hardware in the manner discussed next.

### v4

@@ I Introduction
It would seem, then, that one simply needs to find instances of Eq. (1) whose approximate optimization with the best Monte-Carlo algorithms is computationally hard (i.e., exhibits or at least hints at superpolynomial time complexity) in order to identify good candidates for demonstrating time-complexity quantum advantage in heuristic approximate optimization. However, Ref. [17] provides evidence that short-range rugged-energy-landscape spin glasses may be heuristically approximately optimized classically in quadratic time with approximate contractions of tensor networks, though the wall time for cubic lattices with that heuristic proves to be too large to be practical for large systems. Here we present a major advance over that result: We demonstrate a very simple deterministic tensor-network heuristic that is able to compute in linear time (and space) at least some (but not necessarily all) spin configurations that have energies within a few percent of the ground-state energy for the following three versions of Eq. (1): (1) the cubic-lattice model that is used in the D-Wave approximate optimization experiments in Refs. [5] and [6], (2) a planted-solution, cubic-lattice model whose hardness (i.e., time to exact optimization) for state-of-the-art Monte Carlo heuristics is orders of magnitude more than the model, and (3) a cubic-lattice model that encodes (unweighted) Max-Cut on random 3-regular graphs. Unlike the heuristic in Ref. [17], the present one is able to address very large cubic-lattice systems in a practical amount of time. We also provide an explanation for why this heuristic implies that a quantum algorithm must have sublinear time complexity in order to have quantum advantage in the time complexity of heuristic approximate optimization of classical spin glasses that have correlations that decay with euclidean distance, such as models (1) and (2) just mentioned.
@@ V Summary and Discussion
Another augmentation of the method here could be along the lines of combining it with Monte Carlo similar to what is done in Refs. [44, 45, 46, 29]. In those works the sampling bias from approximate tensor-network contractions is corrected with the Metropolis scheme; in the case of the present method the Metropolis scheme could correct the bias that arises from finite fragment sizes.
We are hopeful that such improvements of the present algorithm may provide a new avenue for gaining insights into the properties of classical spin-glass models studied in condensed matter. For example, questions remain open regarding the nature of the low-temperature phase of the cubic-lattice classical model [47]. Pure Monte Carlo methods have been the state-of-the-art approach for this problem, but the hybrid methods that we propose would operate according to very different principles and may thereby yield new insights. As another example, Ref. [33] points out that the tunability of the tile-planted-solution model can allow for a systematic study of the interplay between disorder and frustration; the algorithms that we propose may provide a complementary route to pure Monte Carlo methods in such a study as well.
We explained that the algorithm may be parallelized through simultaneous use of multiple fragments. This would very substantially reduce the simulation time of our heuristic from what is reported here, which would allow the approximate optimization of much larger systems in a practical amount of time. We also noted that one of the sources of error in our implementation of the algorithm, finite temperature, could be eliminated without changing the time complexity: the ordinary tensors could be replaced by tropical tensors [37]. This would likely reduce the energy errors reported in this work at the same fragment sizes, but would not be compatible with adapting the algorithm to specialized photonic hardware in the manner discussed next.
@@ Acknowledgements.
- Ferris [2015] A. J. Ferris, Unbiased monte carlo for the age of tensor networks, arXiv preprint arXiv:1507.00767 (2015).
- Huggins et al. [2017] W. Huggins, C. D. Freeman, M. Stoudenmire, N. M. Tubman, and K. B. Whaley, Monte Carlo tensor network renormalization, arXiv preprint arXiv:1710.03757 (2017).
- Frías Pérez et al. [2023] M. Frías Pérez, M. Mariën, D. Pérez García, M. C. Bañuls, and S. Iblisdir, Collective Monte Carlo updates through tensor network renormalization, SciPost Physics 14, 123 (2023).